The driver of a 2.0 × 103 kg red car traveling on the highway at 45m/s slams on




his brakes to avoid striking a second yellow car in front of him, which had come




to rest because of blocking ahead as shown in above Fig. After the brakes are




applied, a constant friction force of 7.5 × 103 N acts on the car. Ignore air




resistance.




(a) Determine the least distance should the brakes be applied to avoid a




collision with the other vehicle?




(b) If the distance between the vehicles is initially only 40.0 m, at what speed




would the collision occur?




(c) Write your conclusive observations on the result obtained from this




numerical. I.e. the importance of Physics in daily life.







Solution




3




3




:




2.0 10




45 /








7.5 10




Given data in this Question




mass of the car is kg




velocity of the car is u m s




final velocity v




Friction force is kg




 




 




 




 

1
Expert's answer
2021-12-09T09:14:07-0500

(a)


"d=v^2/(2a)=v^2/(2·F/m)=270\text{ m}."

(b)


"v_f=\sqrt{v^2-2xa}=\sqrt{v^2-2xF/m}=42\text{ km/h}."

(c) Physics gives us the values that will save our lives on roads.